OpenStudy (anonymous):

Resultant capacitor: Is there any method to obtain a specific capacitance from a specific number of capacitors ??

OpenStudy (anonymous):

I mean not by guessing.

OpenStudy (anonymous):

If I nterpret your question correctly, Yes. If you have N capacitors connected in parallel then \[C _{total}= C _{1}+C _{2}+......+C _{N} \] If you have N capacitors connected in series then\[\frac{ 1 }{C _{total} }= \frac{ 1 }{ C _{1} }+\frac{ 1 }{ C _{2} }+....+\frac{ 1 }{ C _{N} }\]

OpenStudy (anonymous):

That is the general law. I mean if I have 4 capacitors of 2 f and want to get: 8f (all para 2f (2 para series with 2 para 1.5f (3series with 1 para 0.5f ( all series by using guessing it's OK and i can solve it but I mean algorithmic way

OpenStudy (anonymous):

there is no guessing involved. I don't understand what you are trying to do.

OpenStudy (anonymous):

when you have 4 capacitors of 4f each and you want to get 1.5f as an equivalent capacitance from them all how can you do this?? I used guessing to say that (3 parallel) and make them series with the forth one to get 1.5f.

OpenStudy (anonymous):

I see.. First, It may not be possible. actually I think it is impossible since the smallest capacitor you can get is four connected in series which gives 1uF. You cannot connect any in parallel because the net Capacitance would be too large.

OpenStudy (anonymous):

oh..uh.... Can just put 2 instead of 4 and all will work sorry for that

OpenStudy (anonymous):

2 farad

OpenStudy (anonymous):

4 capacitors of 2 f to get 1.5f

OpenStudy (anonymous):

You can't put any in parallel with a series array for that would exceed your goal. Since you want 1.5 to get that you could put a 2 uF in series with a 6uF. I got this by assuming that I will need a series connection of two or more caps. Since I need a minimum of two caps I will assume that one is 2 uF. My equation is \[1/1.5 = 1/2 +1/x\] Solving for x I get x = 6 which I can make with the three other 2 uF connected in parallel!!!

OpenStudy (anonymous):

OK ,but you still need a logic to solve this.I hope to get some magical equation to just plug in numbers and get the answer

OpenStudy (anonymous):

There is no single equation for this type of problem. First of all given an arbitrary set of capacitors you will only be able to obtain a limited number of combinations therefore values. If you seek a specific final value from a given set you must start be eliminating certain combinations as I did. Set up an equation with the desired cap on one side and a reasonable starting combination on the other side as I had done. Solve for the unknown capacitance and see if it is achievable with the remaining capacitors. If not then try another starting configuration and try again.

OpenStudy (anonymous):

OK thanks